# ZNorm: Z-Score Gradient Normalization for Deep Neural Networks

### v4

@@ 4 Theoretical Analysis
$d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}=d\left(\frac{1}{d}\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})\right)^{2}=\frac{1}{d}\left(\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})\right)^{2}.$
This expression represents the squared mean of the gradient components scaled by the dimension $d$. Now, by the Cauchy-Schwarz inequality, we have:
@@
$\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})^{2}\geq\frac{1}{d}\left(\sum_{i=1}^{d}\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})\right)^{2}=d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}.$
This inequality ensures that the term $\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}$ is non-negative:
$\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}\geq 0.$
Therefore, the inner product $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ can be expressed as:
@@
$(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})=\frac{\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}-d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}}{\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}}\geq 0.$
For $\hat{g}^{T}g$ to be strictly positive, $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ must not be a constant vector (i.e., not all $\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})$ are the same), ensuring that $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}>0$ and $\|\nabla\mathcal{L}(\mathbf{\theta^{(l)}})\|_{2}^{2}>d\mu_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}^{2}$. This condition holds under typical scenarios in optimization where the gradient vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ varies across its components. The above proof shows that $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ is positive under general conditions, but it may not always be positive in specific cases. Particularly, if all components of the gradient are identical (i.e., $\nabla\mathcal{L}(\mathbf{\theta_{(i)}^{(l)}})=\text{constant}$), the standard deviation $\sigma_{\nabla\mathcal{L}(\mathbf{\theta^{(l)}})}=0$, making the inner product undefined or zero. Therefore, while Z-score standardization generally preserves the descent direction in typical scenarios, it may not always do so in extreme cases. Therefore, under these conditions, $(\Phi_{Z})^{T}\nabla\mathcal{L}(\mathbf{\theta^{(l)}})>0$, which implies that the Z-score normalized gradient $(\Phi_{Z})$ indeed preserves the descent directionality. Despite the potential changes in individual components caused by Z-score standardization, the overall direction of the gradient vector $\nabla\mathcal{L}(\mathbf{\theta^{(l)}})$ that leads to a decrease in the objective function is preserved, proving the theorem.
@@ 5 Experimental Results
5.1 Experimental Setting
All experiments were performed using the Adam optimizer [6], with the baseline referring to the standard Adam gradient. The experiments involved Gradient Clipping [13], Gradient Centralization [22], and ZNorm (Ours). Additionally, Weight Decay 3 and Weight Decay 4 correspond to decay rates of 1E-3 and 1E-4, respectively. For clipping, the value was 0.1. For image classification, we used the CIFAR-10 [7] and PatchCamelyon [20] datasets. All datasets were not augmented.
@@ 5.1 Experimental Setting
5.2 Image Classification Results for CIFAR-10
In this section, we assess the effectiveness of ZNorm on the CIFAR-10 dataset [7] across various neural network architectures. ZNorm consistently improves performance across different CNN architectures, offering better test accuracy and stability compared to other methods.
Table 1 provides detailed experimental results, showing how ZNorm enhances performance across multiple models. For instance, in ResNet-56 [3], ZNorm achieves a test accuracy of 0.812, outperforming the baseline (0.802) as well as methods like Gradient Centralization (0.804) and Gradient Clipping (0.747). Similarly, ResNet-101 [3] shows improved results with ZNorm, reaching a test accuracy of 0.820, surpassing the baseline (0.770) and other techniques.
In DenseNet-121 [4], ZNorm boosts test accuracy to 0.799, outperforming the baseline (0.759) and Gradient Centralization (0.784). DenseNet-169 [4] also benefits from ZNorm, achieving the highest test accuracy of 0.802, compared to the baseline (0.766) and other methods.
Finally, MobileNetV2 [18] with ZNorm reaches a test accuracy of 0.7801, outperforming the baseline (0.763) and other techniques. These results demonstrate ZNorm’s consistent ability to improve accuracy across a diverse set of models and architectures.
@@ 5.2 Image Classification Results for CIFAR-10
[…]
@@ 5.3 Image Classification Results for Breast Tumor
We evaluated the effectiveness of the Z-Score Normalization (ZNorm) method on the PatchCamelyon dataset [20], which consists of breast cancer pathology images used to determine whether the patchified images contain tumors. This real-world dataset allows us to further assess the performance of ZNorm beyond the controlled CIFAR-10 experiments.
We experimented with ResNet-56, ResNet-101, and ResNet-152 models [3]. As shown in Table 2, ZNorm consistently outperformed other normalization techniques across these models. For ResNet-56, ZNorm achieved a test accuracy of 0.915, significantly higher than the baseline accuracy of 0.880 and other methods such as Gradient Centralization (0.887) and Gradient Clipping (0.887). Similarly, for ResNet-101, ZNorm reached the highest test accuracy of 0.917, surpassing both the baseline (0.870) and the other methods. In the case of ResNet-152, ZNorm also outperformed other approaches, achieving a test accuracy of 0.912, compared to the baseline accuracy of 0.884.
These results demonstrate that ZNorm can effectively improve performance on real-world datasets like PatchCamelyon, validating its potential in medical imaging tasks such as tumor prediction.
@@
5.4 Image Segmentation for Brain Tumor
For the brain tumor segmentation task, we utilized the LGG MRI dataset [23], which is aimed at detecting and segmenting tumors in brain MRI images. Three well-established models in the domain were applied: ResNet50-Unet [16], U-net++ [24], and Attention UNet [25].The performance comparison shown in Table 3 highlights that ZNorm consistently outperforms other methods across various models. In the ResNet50-Unet [16], ZNorm achieved the highest test F1 score of 0.917, a Tversky accuracy of 0.914, and the lowest Hausdorff distance of 2.663, outperforming the baseline and other gradient handling techniques like Gradient Centralization and Weight Decay. Similarly, U-net++ [24] produced a strong Tversky accuracy of 0.910 with ZNorm, coupled with competitive results in other metrics, further demonstrating its effectiveness. In the Attention UNet [25], ZNorm recorded the best Tversky accuracy (0.928) and the lowest Hausdorff distance (2.947), while maintaining a high F1 score of 0.894. These findings demonstrate that ZNorm consistently delivers more accurate and precise tumor segmentation compared to other normalization techniques.
The qualitative results in Figure 2 illustrate how ZNorm produces segmentation masks that closely match the ground-truth, with significant improvements over baseline methods, Gradient Centralization [22], Gradient Clipping [13], and different weight decay settings [10]. These results clearly show that ZNorm enhances the robustness and accuracy of brain tumor segmentation across multiple architectures.
@@ 6 Conclusion
We proposed ZNorm, a novel method for gradient normalization, which has proven to significantly enhance the performance of deep neural networks. By normalizing both the mean and variance of the gradients, ZNorm offers consistent gradient scaling across layers, effectively mitigating issues like vanishing and exploding gradients. Our experimental results, spanning across a range of architectures and tasks such as image classification on CIFAR-10 and PatchCamelyon datasets and brain tumor segmentation on the LGG MRI dataset, consistently demonstrate that ZNorm outperforms existing methods like Gradient Centralization and Gradient Clipping. ZNorm leads to better generalization and higher accuracy in both classification and segmentation tasks compared to other normalization techniques. These results highlight ZNorm’s versatility and robustness, making it a valuable tool for improving the performance of deep learning models across various applications and architectures. Given these advantages, ZNorm stands out as a promising method for future work in deep neural network optimization and beyond.

### v5

@@ 3.1 Embedding ZNorm to Adam
$=\delta^{(l)}\cdot\frac{\partial}{\partial\mathbf{\theta}^{(l)}}\left(f(\mathbf{z}^{(l-1)})+\mathbf{z}^{(l-1)}\right)$
@@
$=\delta^{(l)}\cdot\left(\frac{\partial f(\mathbf{z}^{(l-1)})}{\partial\mathbf{\theta}^{(l)}}+\frac{\partial\mathbf{z}^{(l-1)}}{\partial\mathbf{\theta}^{(l)}}\right)$
Applying the chain rule, we obtain:
@@
$\nabla\mathcal{L}(\mathbf{\theta}^{(l)})$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{z}^{(L)}}\cdot\prod_{k=l+1}^{L}\left(\frac{\partial f(\mathbf{z}^{(k-1)})}{\partial\mathbf{z}^{(k-1)}}+1\right)\cdot\frac{\partial\mathbf{z}^{(l)}}{\partial\mathbf{\theta}^{(l)}}$ (10)
$=\delta^{(L)}\cdot\prod_{k=l+1}^{L}\left(\frac{\partial\mathbf{z}^{(k)}}{\partial\mathbf{z}^{(k-1)}}+1\right)\cdot\frac{\partial\mathbf{z}^{(l)}}{\partial\mathbf{\theta}^{(l)}}$
The key observation is that even when $\frac{\partial f(\mathbf{z}^{(l-1)})}{\partial\mathbf{\theta}^{(l)}}$ becomes very small, the additional $1$ term in the product ensures that the gradient $\nabla\mathcal{L}(\mathbf{\theta}^{(l)})$ maintains a non-negligible magnitude [7]. This mechanism effectively prevents gradient vanishing and provides stability to the ZNorm training process. The final gradient expression with skip connections can be formally written as:
$\nabla\mathcal{L}(\mathbf{\theta}^{(l)})=\delta^{(L)}\cdot\prod_{k=l+1}^{L}\left(\frac{\partial\mathbf{z}^{(k)}}{\partial\mathbf{z}^{(k-1)}}+1\right)\cdot\frac{\partial\mathbf{z}^{(l)}}{\partial\mathbf{\theta}^{(l)}}$ (11)
The skip connection effectively keeps the gradient $\nabla\mathcal{L}(\mathbf{\theta}^{(l)})$ from becoming zero or extremely small, thereby improving the training stability of ZNorm and ensuring consistent learning progress.
@@ 4 Experimental Results
[…]
@@
Experimental Settings. All experiments were performed using the Adam optimizer [11], with the baseline referring to the standard Adam gradient. The experiments involved Gradient Clipping [20], Gradient Centralization [32], and ZNorm (Ours). Additionally, Weight Decay 3 and Weight Decay 4 correspond to decay rates of 1E-3 and 1E-4, respectively. For clipping, the value was 0.1. For image classification, we used the CIFAR-10 [12] and PatchCamelyon [30] datasets. All datasets were not augmented. All architectures used in our experiments (ResNet, DenseNet, VGG-16, U-Net variants) include batch normalization layers by default.
Evaluation Metric: Hausdorff Distance. The Hausdorff Distance is a metric used to evaluate the maximum discrepancy between two sets of points, commonly employed in segmentation tasks to measure the spatial alignment between the predicted segmentation mask and the ground truth mask. Specifically, it quantifies the greatest distance from a point in one set to the closest point in the other set. For two sets of points $P$ (predicted mask) and $G$ (ground truth mask), the Hausdorff Distance is defined as:
$H(P,G)=\max\{\sup_{p\in P}\inf_{g\in G}d(p,g),\sup_{g\in G}\inf_{p\in P}d(g,p)\}$ (12)
where $d(p,g)$ is the Euclidean distance between point $p\in P$ and point $g\in G$. The metric calculates the maximum of the minimum distances between the two sets, capturing the largest deviation in their alignment. In segmentation tasks, a lower Hausdorff Distance indicates better spatial alignment between the predicted mask and the ground truth mask, highlighting the precision of the segmentation model in delineating boundaries.
4.1 Image Classification Results for CIFAR-10
In this section, we assess the effectiveness of ZNorm on the CIFAR-10 dataset [12] across various neural network architectures. ZNorm consistently improves performance across different CNN architectures, particularly in skip-connected networks. Table 1 provides detailed experimental results, showing how ZNorm enhances performance across multiple architectures. For instance, in ResNet-56 [7], ZNorm achieves a test accuracy of 0.812, outperforming the baseline (0.802) as well as methods like Gradient Centralization (0.804) and Gradient Clipping (0.747). Similarly, ResNet-101 [7] shows improved results with ZNorm, reaching a test accuracy of 0.820, surpassing the baseline (0.770) and other techniques. In DenseNet-121 [8], ZNorm boosts test accuracy to 0.799, outperforming the baseline (0.759) and Gradient Centralization (0.784). DenseNet-169 [8] also benefits from ZNorm, achieving the highest test accuracy of 0.802, compared to the baseline (0.766). However, in networks without skip connections, such as VGG-16 [27] and Xception [3], ZNorm shows comparable but not superior performance (0.835 vs 0.845 for VGG-16, 0.728 vs 0.751 for Xception). These results demonstrate that ZNorm is particularly effective in improving accuracy for skip-connected architectures, where maintaining stable gradient flow is crucial for performance.
@@ 4.2 Image Classification Results for Breast Tumor
We tested ZNorm on the PatchCamelyon dataset [30], which contains breast cancer images, to check if it works well on real medical data. Using three different ResNet models (ResNet-56, -101, and -152) [7], our results in Table 2 show that ZNorm works better than other methods. ZNorm achieved high test accuracy across all models: 0.915 with ResNet-56 (compared to baseline 0.880), 0.917 with ResNet-101 (baseline 0.870), and 0.912 with ResNet-152 (baseline 0.884). These results show that ZNorm works well for real medical tasks like finding tumors in breast cancer images.
4.3 Image Segmentation for Brain Tumor
For the brain tumor segmentation task, we tested the LGG MRI dataset [35] using three skip-connected models: ResNet50-Unet [23], U-net++[36], and Attention U-Net[37]. These models use skip connections, which help ZNorm work better as explained earlier. Looking at Table 3, ZNorm does better than other methods across all models. Using ResNet50-Unet, ZNorm got the best test F1 score (0.917) and Tversky score (0.914), and the smallest Hausdorff distance (2.663). For U-net++, while Gradient Centralization had a slightly better F1 score (0.901), ZNorm still did well (F1: 0.898, Tversky: 0.910). With Attention U-Net, ZNorm got the best Tversky score (0.928) and smallest Hausdorff distance (2.947). These results show that ZNorm makes tumor segmentation better across different models, especially by reducing Hausdorff distance while keeping good F1 and Tversky scores. Looking at the mask in Figure 2, we can see that ZNorm makes tumor outlines that match the real tumors better than other methods like Gradient Centralization [32], Gradient Clipping [20], and different weight decay settings [17].
@@ 4.3 Image Segmentation for Brain Tumor
5 Discussion
5.1 Impact on Skip-Connected Architectures.
The role of skip connections in ensuring the stability of ZNorm was particularly evident in our results. By preserving gradient flow and preventing vanishing gradients, skip connections in architectures like ResNet and U-Net enhanced ZNorm’s ability to stabilize training. This synergy between ZNorm and skip connections highlights its suitability for these architectures, making it a valuable tool for efficiently training networks that leverage residual and dense connections.
5.2 Practical Implications
ZNorm’s simplicity and ease of implementation make it a practical choice for real-world applications. Unlike normalization methods that require architectural modifications or additional hyperparameters, ZNorm operates directly on gradients, allowing it to be seamlessly integrated into existing optimization pipelines. This adaptability reduces the computational and developmental overhead typically associated with implementing new training methods, making ZNorm particularly appealing for applications requiring rapid prototyping and deployment.
5.3 Limitations and Future Research
While our study demonstrates the effectiveness of ZNorm across various tasks and architectures, it also highlights several directions for extended research. One limitation is that the experiments were primarily conducted on relatively small- to medium-scale datasets such as CIFAR-10 and medical datasets like PatchCamelyon and LGG MRI. These datasets effectively showcase the benefits of ZNorm, but future research will focus on larger and more diverse datasets, such as ImageNet, to validate its scalability and generalizability. Conducting experiments on ImageNet will provide critical insights into whether the performance improvements observed in this study remain consistent in larger-scale and more complex tasks.
Another area for exploration involves addressing the challenges ZNorm may encounter near convergence. As gradient magnitudes become extremely small during this phase, there is a risk of instability in the normalization process. Although the use of skip connections mitigates this issue to some extent, future work will explore adaptive mechanisms to further stabilize ZNorm, ensuring robustness across all stages of training.
These extensions will build upon the findings of this study, offering opportunities to further enhance the training efficiency, stability, and scalability of ZNorm across an even broader range of applications and domains. By addressing these areas, we aim to establish ZNorm as a foundational tool for efficient deep learning in increasingly complex and diverse settings
@@ 6 Conclusion
In this paper, we proposed Z-score Gradient Normalization (ZNorm), a novel and efficient method for training skip-connected neural networks. By normalizing both the mean and variance of gradients, ZNorm mitigates common challenges such as vanishing and exploding gradients without requiring architectural modifications. This simplicity ensures seamless integration into existing training pipelines while enhancing the training stability and performance of deep neural networks. Our experimental results demonstrate the effectiveness of ZNorm across diverse tasks and datasets. On image classification tasks with CIFAR-10 and PatchCamelyon, ZNorm consistently outperformed baseline methods and other gradient adjustment techniques. In medical image segmentation, ZNorm achieved superior results on the LGG MRI dataset, particularly in reducing the Hausdorff Distance while maintaining high Tversky and F1 scores. These results highlight ZNorm’s capability to improve performance and generalization in both standard and high-impact applications, especially in skip-connected architectures like ResNet and U-Net. In conclusion, ZNorm offers a robust and straightforward approach to optimizing deep neural network training. Its ability to enhance performance without modifying network architecture positions it as a valuable tool for advancing efficient training methods in deep learning. Future work will build upon these findings to refine and expand ZNorm’s applicability in the evolving landscape of deep learning research.